OpenStudy (anonymous):

hhhhhhhhhhhhhhhhhhhhhhhhhheeeeeeeeeeeeeeeeeeeeeeeellllllllllllllllllllllllllllllllllllllllllpppppppppppppppppppppppppppppppppppppppppp

OpenStudy (anonymous):

What do you need help with?? :D

OpenStudy (anonymous):

Part A) circumference = PI * diameter PI = circumference / diameter PI = 21.98 / 7 PI = 3.14 PI = 18.84/6 PI = 3.14

OpenStudy (anonymous):

Part B) Area = PI * diameter^2 / 4 PI = 4 * Area / diam^2 PI = 4 * 38.465 / 7*7 PI = 153.86 / 49 PI = 3.14 Circle B PI = 4 * Area / diam^2 PI = 4 * 28.26 / 6 * 6 PI = 113.04 / 36 PI = 3.14

OpenStudy (anonymous):

so that should get me my answer

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

i hope

OpenStudy (anonymous):

kk thanks

OpenStudy (anonymous):

Part C) Conclusion - PI = 3.14 and is the same for both circles and is the same for EVERY circle.

OpenStudy (anonymous):

can i use that as my explanation

OpenStudy (anonymous):

????????????

OpenStudy (anonymous):

Sure

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

np ^_^

OpenStudy (anonymous):

lol